OpenStudy (anonymous):

I can't figure out this: 289+196-2(17)(14) cos50 I am supposed to get 179.0331, but I do not know what I am doing wrong.

OpenStudy (anonymous):

whats the original problem

OpenStudy (anonymous):

or is that it

OpenStudy (anonymous):

Thats the equation that I plugged in, just like the example on the worksheet

OpenStudy (mertsj):

c^2=289+196-2(17)(14) cos50 c^2=485-476 cos 50 c^2=485-305.9669 c^2=179.0331 c=13.3803

OpenStudy (anonymous):

Yay thank you!